OpenStudy (anonymous):

A spaceship moving with an initial velocity of 58.0 meters/second experiences a uniform acceleration and attains a final velocity of 153 meters/second. What distance has the spaceship covered after 12.0 seconds?

OpenStudy (anonymous):

IS IT AN ALIEN SPACESHIP?

OpenStudy (welshfella):

use one of the equations of motion for uniform acceleration

OpenStudy (welshfella):

u = 58 , v = 153, t = 12

OpenStudy (anonymous):

A.6.96 × 102 meters B.1.27 × 103 meters C.5.70 × 102 meters D.1.26 × 102 meters E.6.28 × 102 meters

OpenStudy (welshfella):

d = (u +v) * t ----- 2

OpenStudy (anonymous):

k figured it out its B

OpenStudy (welshfella):

yes that's correct

OpenStudy (perl):

$$ \Large \rm { \\ Definition:\\ d= \overline v \cdot t \\~\\ For ~ uniform ~ acceleration ~it~ is ~ true~ that \\ \overline v = \frac{v_i + v_f}{2} \\~\\ \text{Therefore }\\~\\ \\ d =( \frac{v_i + v_f}{2})\cdot t }$$